OpenStudy (anonymous):

Find all the zeros for the equation -3x^2+27x^2+1200=0

OpenStudy (anonymous):

-3x^2 + 27x^2 + 1200 = 0 Start by combining the x^2 terms. 24x^2 + 1200 = 0 Then divide everything by 24...

OpenStudy (anonymous):

@vivianrush are you working on this one still?

OpenStudy (anonymous):

i divide 1200 by 24 and got 50

OpenStudy (anonymous):

You also divide the 24x^2 term by 24 to get just x^2. It should look like: (24/24)x^2 + 1200/24 = 0 x^2 + 50 = 0

OpenStudy (anonymous):

so then its x^2 +50=0 im suppose to find all the zeros of the equation

OpenStudy (anonymous):

Can they be imaginary number roots? Or are you only allowed to have real number roots in this class?

OpenStudy (anonymous):

real number roots

OpenStudy (anonymous):

x^2 + 50 = 0 Subtract 50 from both sides: x^2 + 50 - 50 = 0 - 50 x^2 = -50 Then you would need to take the square root of both sides... ordinarily you would get a "plus or minus something" as your two zeroes... but in this case, the stuff under the square root is a negative 50... you can't take the square root of a negative number and get real number roots (or "zeroes").

OpenStudy (anonymous):

thank you so much. i understand this a lot better now

OpenStudy (anonymous):

If it had been like x^2 = 25 or whatever, then you would take square root and get zeroes of x equal to plus and minus 5 (or "5 and -5"). But if the stuff under the square root is negative, the answer is "no real zeroes" or something similar.

OpenStudy (anonymous):

Glad to help :)